OpenStudy (anonymous):

For the function f(x) = 2cosx, find all numbers c in the interval [0, pi/2] that satisfy the Mean Value Theorem.

OpenStudy (anonymous):

the derivative is \(-2\sin(x)\) so solve \[-2\sin(x)=\frac{2\cos(\frac{\pi}{2})-2\cos(0)}{\frac{\pi}{2}}\]

OpenStudy (anonymous):

I did that. I got c=1, but it's not working out.

OpenStudy (anonymous):

the right hand sides is \[\frac{-2}{\frac{\pi}{2}}=-\frac{4}{\pi}\]

OpenStudy (anonymous):

so somehow you are supposed to solve \[-2\sin(c)=\frac{-4}{\pi}\] or \[\sin(c)=\frac{2}{\pi}\] you will need to take the arcsine of both sides, that is the best you can do

OpenStudy (anonymous):

Oh, thanks. That's what I got the first time, but it looked so ugly!

OpenStudy (anonymous):

<3